OpenStudy (anonymous):

**Will Give Medals** A piece of machinery is worth $1600. The rate of depreciation of the value V (in dollars) at time t (in years) is proportional to the square root of its value V . The piece of machinery will be worth $900 two years later. When will the piece of machinery be worth $625?

OpenStudy (anonymous):

i need to make a differential equation which will lead into integrals

OpenStudy (amistre64):

dv/dt = sqrt(v) dv/sqrt(v) = dt 1/2 sqrt(v) = t + C sqrt(v) = 2t + C v = (2t + C)^2 ............................. v' = 2(2t + C) = 2 sqrt(v) just taking a stab at it

OpenStudy (amistre64):

C = 40 since 40^2 = 1600

OpenStudy (amistre64):

yeah, this function doesnt decrease over time does it

OpenStudy (amistre64):

and your answer key gives us? if you have one ...

OpenStudy (shiraz14):

@amistre64: dv/dt = k*sqrt(v) (where k is some arb. const) rather than dv/dt = sqrt(v) (we can't assume that k=1 in this context).

OpenStudy (shiraz14):

@bpaquin007 & @amistre64: So the equation written by amistre64 becomes modified to read as follows: dv/dt = k*(√v) ∫ dv/k*(√v) = ∫ dt (2/k)*(√v) = t + C V = [k*(t+C)/2]^2 ---------- (1) Given that when t=0, V=1600 & when t=2, V=900, we have: kC/2 = 40 ----------- (2) k*(2+C)/2 = 30 ------- (3) From (2), we have: kC = 80 ------------- (4) From (3), we have: 2k + kC = 60 --------- (5) Substituting (4) into (5), we have: 2k = -20 k = -10 Substituting k = -10 into (4), we have: C = -8 Substituting these values of k & C into (1), we have: V = [(-5)*(t-8)]^2 ---------- (6) Given that V = 625, we now have from (6): 625 = [(-5)*(t-8)]^2 25 = (-5)*(t-8) t-8 = -5 t = 3 Hence, the machinery will be valued at $625 after 3 years from its initial valuation.

OpenStudy (amistre64):

thats good, i was thinking the constant of proportionality, k, was going to be taken care of in the arbitrary constant .... its good to know better now :)